A man is 41 years old and his son is 9. In how many years will the father be three times as old as the son?: A man is 41 years old and his son is 9. In how many years will the father be three times as old as the son?
Answer by nerdybill(1052)

.
Let x = number of years until father will be 3 times as old as son
.
From the statement of the problem, we get our equation:
41+x = 3(9+x)
41 + x = 27 + 3x
41 = 27 + 2x
14 = 2x
7 = x
.
Therefore, in 7 years the man will be three times as old as the son.